Marks for this submission: 1.00/1.00 The national average SAT score (for Verbal and Math) is 1028 with a standard deviation of 92. What is the probability that a randomly selected score exceeds 1200? ete o 1.00 Select one: estion a. 0.41 b. 0.88 c. 0.59 d. 0.12 e. Cannot be determined Check In the university library elevator there is a sign indicating a 16-person limit as well as a weight limit of 25 weight of stude 27 lbs. Reference: Ref 5-17 00 lbs. Suppose that the nts, faculty, and staff is approximately Normally distributed with a mean weight of 150 ibs and a standard deviation of plete t of 1.00 question What is the probability that the random sample of 16 people in the elevator will exceed the weight limit? Round your answer to 4 decimal places. Answer: Check The distribution of total body protein in adult men with liver cirrhosis is Normal with mean 9.8 kg and standard deviation 0.1 kg. A random sample of 25 adult men with liver cirrhosis is taken. What percentage of men in the population have a total body protein that on 11 mpieteQuestion 12 Not complete Points out of 1.00 P Flag question level of nitrogen oxides (NOX) in the exhaust of cars of a particular model varies with a distribution with mean 0.25 grams per mile (g/mi) and standard deviation 0.05 g/mi. Government regulations call for NOX emissions no higher than 0.3 g/mi. A company has a sample of 44 cars of this model in its fleet. What is the probability that the average NOX level of these cars is above the 0.3 g/mi limit? Select one: a. 0.2525 b. 0.1586 c. 1.7 10A-11 d. 0.0228 O e. Cannot be determined Check The average monthly mortgage payment including principal and inte deviation is $180, there is a 75% chance that the mean mortgage payment of a random sample of 16 loans is more than what? Question 13 rest is $982 in the United States. If the population standard Not complete Ptsu of 100Select one: P Flag question a. 860.65 o b. 951.65 c. 1103.41 d. 1012.35 e. Cannot be determined Checkcould you solve these 4 questions. Thank you
Answer #1

1) The national average SAT score is 1028 and standard deviation is 92. Calculate the probability that the randomly selected score exceeds 1200. P(X> 1200) = 1-P(X 1200) x-u , 1200-μ 1200-1028 92 1-P(Z -1-0.97 -0.03 1.89) 2) The sample size is 16. The mean is 150 and standard deviation is 27 Now the mean of the total sample is 2400 (16x150)and standard deviation of the total sample is 108 (4x28) Therefore,P(x>2500)-1-P(X, S 2500) 2500- 2400 -1- P(Z s0.93) = 1-0.8228 = 0.1772 3) Calculate the require probability X>0.3)=1-P( X 0.3 =1-Plz 03-025 1- PZS 6.63) 1.710-11 0.054) That is, 951.65 Since, P( X > 0.75 7 17 1-P(78 z) 0.75 P(T& z) = 0.25 From the standard z table values, observe that the critical value of z for the two tail test at the left tail test and the corresponding cumulative area of 0.25 is -0.674 Therefore, z =-0.674 =-0.674 7isa-06741 180 V44 = 982-0.674 982-30-35 951 .65